OpenStudy (anonymous):

What will be the impact of collision on two vehicles moving at opposite directions at the same acceleration rate if one is slightly heavier than the other

OpenStudy (anonymous):

I'm not sure what you are asking can you post what problem you are working on?

OpenStudy (anonymous):

NetForce = a(m1 - m2) Be careful on your use of "heavier", which implies weight instead of what should be designated as larger mass.

OpenStudy (anonymous):

The acceleration is immaterial. All you need to know is the velocity of the two vehicles at the instant of collision. Apply conservation of linear momentum. Then you have an equation in two unknowns( The final velocities of the cars). Then you can solve the problem by applying the coefficient of restitution. And by the way the soltuion given by cbott7 is absolutely, ridiculously wrong

OpenStudy (jh3power):

F=dp/dt=d(mv)/dt=ma ... p=mv ... I=Ft=(delta)mv